# Active-Learning Inspired Ab Initio Theory-Experiment Loop Approach for Management of Material Defects: Application to Superconducting Qubits

### v2

@@ Abstract
Introduction. The rational design of materials has played a pivotal role in advancing various technologies by accelerating the development of materials with tailored properties. This approach has been extensively applied in fields ranging from fuel cell catalysis [1], energy storage materials [2], and photonic structures [3] to high performance polymers [4]. These advances have been enabled by computational tools, high-throughput screening, and increasingly, machine learning, which allow for the optimization of properties including bandgaps [5], thermal conductivity [6], and electrochemical activity [7]. Despite these advances, the vast majority of rational design efforts have focused on bulk properties of materials, with relatively little attention paid to defect energetics [8].
@@
While much of rational materials design has focused on tuning bulk properties, defect energetics play a critical role in determining how atoms migrate, especially across thin films and interfaces. This is particularly important in the design of materials for quantum computing [9, 10], specifically, in superconducting quantum devices, where Josephson junctions (JJs) are formed by sandwiching thin insulating or oxidized layers between superconducting electrodes. In such systems, oxygen diffusion from adjacent capping layers or the ambient environment can lead to the formation of unwanted oxides at the interface, most notably niobium oxide [11, 12]. These oxides host two-level systems (TLSs) that couple to the electric field of the qubit, acting as parasitic resonators that degrade coherence times and limit device performance [13, 14]. Suppressing the formation of these oxides is therefore a key challenge in improving qubit performance. Various strategies in this regard have been explored, with some notable methods such as high-temperature heat treatment and etching seen to reduce TLS losses [11, 15, 16, 17]. One such method, namely encapsulation of the superconducting metal with a barrier surface, has also shown experimental promise [18]. However, such studies have been mostly experimental, with insufficient theoretical investigation exemplifying use of the Materials Genome Loop.
@@
In this work, we introduce a predictive framework that leverages defect energies calculated from first principles to evaluate the formation of niobium oxide in superconducting qubit JJs. By quantifying key microscopic energies—specifically, vacancy formation energies in metal oxides and oxygen interstitial energies in metals, we construct a logistic regression model that captures the likelihood of oxide formation across a range of candidate metals. The result is a phase-space map that distinguishes between effective and ineffective oxygen barrier materials and enables us to correctly predict new successful barrier materials. More broadly, this work demonstrates how rational defect engineering can be used to design atomically coherent, oxidation-resistant interfaces for quantum devices.
Computational details. All ab initio density–functional theory (DFT) calculations were performed with the open-source plane-wave code JDFTx [19, 20]. Ultrasoft pseudopotentials from the GBRV library [21] were employed together with the Perdew–Burke–Ernzerhof functional revised for solids (PBEsol) [22]. Calculations used the generalized-gradient approximation (GGA) and periodic boundary conditions in all directions. Plane-wave kinetic-energy cut-offs were set to 20 hartree for Kohn–Sham wavefunctions and 100 hartree for the charge density.
For each metal-oxide composition we selected the structure with the lowest formation energy reported in the Materials Project database [23]. Supercells were then constructed so that every defect was separated from its periodic images by at least three metal–metal spacings in all Cartesian directions, suppressing spurious defect–defect interactions. The same supercells were used for pristine and defect calculations to ensure strict numerical consistency.
Brillouin-zone integrations employed Monkhorst–Pack $\mathbf{k}$-point meshes, offset by $(0.5,0.5,0.5)$, that were individually converged until the total-energy change was below $10^{-5}\,\text{hartree}$ ($\approx 0.27\,\text{meV}$). Electronic self-consistency cycles were deemed converged when the total-energy difference between successive iterations fell beneath the same $10^{-5}\,\text{hartree}$ threshold.
[…]
@@
where $E(\mathrm{M}_{Nn}\mathrm{O}_{Nm-1})$ is the energy of a supercell with $N$ unit cells of the oxide $\mathrm{M}_{n}\mathrm{O}_{m}$ minus one O atom, and $E(\mathrm{M}_{Nn}\mathrm{O}_{Nm})$ is the energy of a supercell with $N$ unit cells of the oxide $\mathrm{M}_{n}\mathrm{O}_{m}$.
@@
To reinforce this conclusion, we estimated the characteristic diffusion time for an oxygen atom to traverse a 5 nm capping layer. To estimate the diffusion constants, we use the standard Arrhenius expression:
$D=\frac{1}{6}a^{2}\nu\,e^{-\tfrac{E_{d}}{k_{B}T}}$ (8)
where $a\sim 0.3$ nm is the typical jump distance, $\nu\sim 10^{13}$ Hz is a characteristic phonon frequency, and $E_{d}$ is the diffusion barrier energy. Then, using $L\sim 5$ nm as the barrier thickness, we can estimate the corresponding diffusion times:
$\tau=\frac{L^{2}}{2D}$ (9)
For a moderate barrier of $E_{d}=0.5$ eV at room temperature, we find $\tau\sim 40$ ms. Even for higher barriers of $E_{d}\sim 0.8$ eV, the timescale remains on the order of a few minutes. This analysis shows that oxygen diffusion is generally fast enough to occur within experimental time frames across nearly all of our metals, except for those with the very highest barriers. Yet, we also find examples of low-barrier materials that successfully prevent oxide formation, and, conversely, our material with a $\sim$1.5 eV barrier (a $\sim 10^{12}$ increase in timescale over 0.8 eV) still allows for oxide formation. This indicates that oxygen can access alternative pathways (e.g., grain boundaries) to traverse the barrier. All together, these observations suggest that the presence or absence of niobium oxide is not controlled primarily by kinetic limitations, but rather more so by the thermodynamic favorability of the oxygen transfer processes described above. We therefore use only the interstitial and vacancy formation energies as descriptors in our model.
Phase diagram.
Figure 2: Oxide vacancy energy versus metal interstitial energy for a set of metals. Darker green regions indicate increasingly favorable thermodynamic barriers to oxygen diffusion
@@
Figure 2 shows a phase diagram of oxygen vacancy energy versus metal interstitial energy for various candidate metals. Our initial processes in Equations 3, 4, and 5 represent the sequential microscopic steps necessary to form niobium oxide. For a metal to be a good diffusion barrier, we would like each of these processes to be thermodynamically unfavorable ($\Delta E>0$). Each of these conditions corresponds to a linear inequality in our two-dimensional descriptor space, forming a half-plane boundary. In Figure 2, the shaded green regions indicate the zones where each individual process is suppressed. Where these regions overlap, the shading becomes darker, indicating that multiple (or all) oxygen transfer pathways are energetically blocked. Metals located in these darker regions are therefore expected to serve as more effective diffusion barriers.
@@
Most metals explored lie above the diagonal boundary corresponding to Equation 3, which prevents oxygen from leaching out of the surface oxide and entering the barrier metal. The vertical boundary near $E(\mathrm{O}_{I}(\mathrm{M}))\sim-3.5$ eV (from Equation 4) excludes metals that would readily transfer interstitial oxygen into metallic niobium as an interstitial. A stricter boundary near $E(\mathrm{O}_{I}(\mathrm{M}))\sim-6$ eV (from Equation 5) further excludes metals that would facilitate oxygen interstitial transfer into pre-existing niobium oxide vacancies. Both of these vertical cuts suggest that moving leftward in the plot corresponds to increasing thermodynamic resistance to oxide formation. Together, these three conditions indicate that materials in the upper-left quadrant of the phase space are most favorable, while those in the lower-right are least favorable.
At this level of analysis, of the potential barrier materials considered so far, one in particular, zirconium, stands out as satisfying all three criteria and therefore is a particularly promising candidate as a barrier material. Moreover, even if the zirconium is fully oxidized, the combination of Equations 3 and 4 ensures that it will maintain a sharp interface with niobium and thus zirconium oxide is also a particularly promising candidate as an insulating layer for Josephson junctions.
@@
Our initial experimental data confirmed the broad trends predicted by the descriptor phase space, with successful barrier metals appearing toward the upper left of the phase space and unsuccessful metals appearing to the lower right. We then used this data to train the first iteration of our logistic regression model as described above in Logistic regression. Figure 3a shows the results, where experimental outcomes for Al, Ta, TiW3, Pd, Pt, and Au are marked by green or red circles outlining the data points, with green and red indicating whether niobium oxide formation was observed to be suppressed or not. These data were then used to train an initial logistic regression model, which outputs a continuous probability of oxide formation as a function of defect energies. The predicted probabilities are encoded in the figure as the color of the inner solid circles, ranging from green ($P=0$) to red ($P=1$).
Metals were then prioritized for further testing based on their predicted probability of success. After each new experiment, the result was fed back into the model, which was retrained to improve its predictive accuracy. This created a closed theory-experiment loop that progressed over several iterations. By the most recent iteration, shown in Figure 3b, the model had successfully predicted the behavior of five new potential barrier materials: Zr, Hf, Ti, Mo, and Ir, with one unsuccessful case: W. The final model shows strong consistency with experiment and captures smooth, interpretable probability gradients across the descriptor space.
@@
To test whether the per oxygen oxide formation energy is sufficient to explain our logistic regression predictions, we next project a line of constant oxide formation energy (determined by logistical regression holding the slope of the line fixed) back onto the original descriptor phase space, as seen in Figure 3b. The resulting line separates the data just as well as the decision boundary learned by our original logistic regression model, indicating that per-oxygen oxide formation energy alone is sufficient to reproduce the model’s classification performance.
Figure 5: Per-oxygen oxide formation energy of various experimentally tested oxides, colored according to their success (green) or failure (red) to prevent oxidation of the niobium layer
@@
Figure 6: Oxide formation energy versus lattice mismatch percentage against niobium: Metals have been colored according to their logistic predictions, along with outer circles indicating experimental results denoted by their color
@@
Conclusion. In this work, we presented a data-driven framework that integrates DFT defect energetics with logistic regression to design capping layers suppressing niobium oxides in superconducting devices. Two physically meaningful descriptors—oxygen interstitial energies in metals and oxygen-vacancy energies in their oxides—define a learned decision boundary that separates effective from ineffective barriers, prospectively identifying Zr, Hf, and Ta, which we then verified experimentally. Adding lattice mismatch as a secondary screening criterion prioritizes coherent interfaces. Within this framework, Zr and Ta combine low oxide propensity with small mismatch, Hf proves effective albeit more mismatched, and Sc emerges as a new potential candidate.
This work highlights a generalizable methodology for integrating defect-level energetics into the rational design of functional materials. More broadly, the strategy is applicable whenever readily computable, physically controlling descriptors can be identified. By unifying such descriptors with limited experimental studies through carefully selected machine learning techniques, as demonstrated here, we establish a robust closed-loop materials-by-design framework that accelerates the discovery of next-generation materials.
Acknowledgments. This work was supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0706. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force. This work made use of the Cornell Center for Materials Research shared instrumentation facility for the XPS analysis system.
Data availability. The data that support the findings of this article are openly available [28].

### v3

@@ Abstract
Introduction. The rational design of materials has played a pivotal role in advancing various technologies by accelerating the development of materials with tailored properties. This approach has been extensively applied in fields ranging from fuel cell catalysis [1], energy storage materials [2], and photonic structures [3] to high performance polymers [4]. These advances have been enabled by computational tools, high-throughput screening, and increasingly, machine learning, which allow for the optimization of bulk properties including bandgaps [5], thermal conductivity [6], and electrochemical activity [7]. As device dimensions shrink and interfaces dominate performance, local atomic-scale processes become as critical as bulk properties. Defect energetics are therefore widely studied using both first-principles and data-driven approaches [8, 9, 10, 11, 12, 13, 14], yet they are rarely employed as explicit forward-design variables in theory–experiment closed-loop frameworks linked to device-level failure mechanisms.
@@
This design perspective is particularly critical in quantum computing [15, 16], specifically in superconducting quantum devices, where Josephson junctions (JJs) are formed by sandwiching thin insulating or oxidized layers between superconducting electrodes. In such systems, oxygen diffusion from adjacent capping layers or the ambient environment can lead to the formation of unwanted oxides at the interfaces, most notably niobium oxide in niobium-based superconducting JJs [17, 18]. These oxides host two-level systems (TLSs) that couple to the electric field of the qubit, acting as parasitic resonators that degrade coherence times and limit device performance [19, 20].
Suppressing the formation of such oxides is therefore a key challenge in improving Nb JJ qubit performance. Various strategies in this regard have been explored, including high-temperature heat treatment and etching, which reduce TLS losses [17, 21, 22, 23]. Also promising from a materials-design standpoint is encapsulation of the superconducting metal with an oxygen barrier surface, which has demonstrated encouraging experimental performance [24]. However, barrier-material selection and optimization remain largely empirical, with limited predictive guidance.
Figure 1: Schematic overview of the theory–experiment active-learning workflow used in this work.
Within the Materials Genome framework, enabling predictive design of oxidation-resistant barriers naturally invites active-learning–inspired materials-selection strategies, including surrogate optimization [25, 11, 26, 27, 28, 29] and feasibility-based screening [30, 31]. However, such approaches are most commonly deployed in high-throughput, data-rich settings and offer limited physical interpretability, motivating the development of interpretable active-learning frameworks suited to sparse-data materials-design problems.
@@
In this work, we introduce a predictive framework that leverages defect energies calculated from first principles to guide the selection of oxidation-resistant barrier materials for niobium-based superconducting qubits. By quantifying key microscopic energetics, vacancy formation energies in metal oxides and oxygen interstitial energies in metals, we construct an interpretable logistic-regression model that captures the likelihood of interfacial oxide formation across candidate barriers. The resulting descriptor-space map distinguishes effective from ineffective oxygen barriers and enables prediction of new viable materials. The model operates in a closed-loop workflow in which first principles energetics define the initial predictions and guide early experiments. Experimental outcomes then train the model, which helps direct targeted DFT calculations to update the descriptor space, enabling subsequent cycles of experiments, DFT calculations, and model refinement (Fig. 1). More broadly, this framework illustrates how theory and experiment can be integrated in sparse-data settings through an interpretable, active-learning–inspired implementation of the Materials Genome design loop.
Computational details. All ab initio density–functional theory (DFT) calculations were performed using the open-source plane-wave code JDFTx [32, 33] with ultrasoft GBRV pseudopotentials [34] and the PBEsol exchange-correlation functional [35]. Computational parameters, structural models, convergence criteria, and defect-supercell specifications are provided in the Supplemental Material [36] (see also reference [37] therein).
Experimental details. Niobium thin films were deposited on high-resistivity Si(100) substrates by Ar magnetron sputtering and patterned using standard photolithography and reactive-ion etching. Samples received post-fabrication surface treatments prior to characterization. X-ray photoelectron spectroscopy (XPS) measurements were performed to quantify interfacial oxide formation and to classify diffusion-barrier performance. Fabrication protocols, surface treatments, measurement conditions, and spectral fitting procedures are provided in the Supplemental Material [36] (see also references [38, 39, 40] therein).
@@
where $E(\mathrm{M}_{Nn}\mathrm{O}_{Nm-1})$ is the energy of a supercell with $N$ unit cells of the oxide $\mathrm{M}_{n}\mathrm{O}_{m}$ minus one O atom, and $E(\mathrm{M}_{Nn}\mathrm{O}_{Nm})$ is the energy of a supercell with $N$ unit cells of the oxide $\mathrm{M}_{n}\mathrm{O}_{m}$. Because our objective is to construct comparative descriptors rather than reproduce experimental defect formation energies, we evaluate neutral defect energetics in the expressions above for consistency and transferability. Charged-defect calculations require specification of an electron chemical potential that depends sensitively on interface and device conditions and thus introduce ambiguities. Although defect charge states and operating environments may vary, neutral energetics remain positively correlated with charged-defect energetics and therefore serve as informative, transferable descriptors for diffusion-barrier performance.
@@
To reinforce this conclusion, we estimated oxygen diffusion times across a 5-nm capping layer using Arrhenius diffusion models (see Supplemental Material [36]). Traversal occurs on sub-second to minute timescales for representative barriers (0.5–0.8 eV), indicating that oxygen transport is generally rapid within experimental timeframes. However, Figure 2 shows that low-barrier materials can still suppress oxide formation, whereas the highest-barrier case ($\sim$1.5 eV) permits oxidation despite an estimated diffusion time $\sim 10^{12}$ times longer than for a 0.8 eV barrier. This implies that oxygen can access alternative pathways (e.g., grain boundaries) to traverse the barrier. Combined with the lack of correlation between diffusion barriers and observed oxide suppression, these findings indicate primarily thermodynamic rather than kinetic control of oxide formation. We therefore use only the interstitial and vacancy formation energies as descriptors in our model.
Descriptor map. Figure 3 shows a descriptor map of oxygen vacancy energy versus metal interstitial energy for the initial set of candidate barrier metals. This materials set was selected to span established junction constituents and technologically relevant barrier chemistries, including noble metals resistant to oxidation (Au, Pt, Pd); Al, the widely used Al2O3 tunnel-barrier constituent in Nb-based junctions [42]; refractory and Nb-adjacent transition metals (Ta, Ti, Zr, W, Mo), selected for chemical similarity, lattice compatibility, and strong oxygen affinity; and TiWx, a refractory diffusion-barrier alloy widely employed in thin-film metallization stacks [43].
@@
Our initial processes in Equations 3, 4, and 5 represent the sequential microscopic steps necessary to form niobium oxide. For a metal to be a good diffusion barrier, we would like each of these processes to be thermodynamically unfavorable ($\Delta E>0$). Each of these conditions corresponds to a linear inequality in our two-dimensional descriptor space, forming a half-plane boundary. In Figure 3, the shaded green regions indicate the zones where each individual process is suppressed. Where these regions overlap, the shading becomes darker, indicating that multiple (or all) oxygen transfer pathways are energetically blocked. Metals located in these darker regions are therefore expected to serve as more effective diffusion barriers.
Figure 3: Oxide vacancy energy versus metal interstitial energy for a set of metals. Darker green regions indicate increasingly favorable thermodynamic barriers to oxygen diffusion.
@@
Most metals explored lie above the diagonal boundary corresponding to Equation 3, which prevents oxygen from leaching out of the surface oxide and entering the barrier metal. The vertical boundary near $E(\mathrm{O}_{I}(\mathrm{M}))\sim-3.5$ eV (from Equation 4) excludes metals that would readily transfer interstitial oxygen into metallic niobium as an interstitial. A stricter boundary near $E(\mathrm{O}_{I}(\mathrm{M}))\sim-6$ eV (from Equation 5) further excludes metals that would facilitate oxygen interstitial transfer into pre-existing niobium oxide vacancies. Both of these vertical cuts suggest that moving leftward in the plot corresponds to increasing thermodynamic resistance to oxide formation. Together, these three conditions indicate that materials in the upper-left quadrant of the descriptor space are most favorable, while those in the lower-right are least favorable.
At this level of analysis, zirconium stands out as satisfying all three criteria and therefore emerges as a particularly promising barrier material. Even if fully oxidized, Eqs. 3 and 4 indicate that zirconium should maintain a chemically sharp interface with niobium, making zirconium oxide a compelling insulating candidate for Josephson junctions. This thermodynamic picture is consistent with prior studies from the superconducting radio-frequency community documenting zirconium surface segregation and preferential oxidation in niobium–zirconium alloys [44, 45, 46, 47], as well as direct observations of passivating ZrO2 formation on niobium surfaces [48]. In contrast, the present work addresses the defect energetics and interfacial stability governing diffusion-barrier behavior. Building on the thermodynamic predictions presented here, recent experimental studies of Nb/ZrOx/Nb junctions have reported exceptionally abrupt interfaces, consistent with the predicted oxygen segregation behavior [49].
@@
Our initial experimental dataset corresponds to the circled points in Fig. 4a, selected as a representative subset of the ab initio set to broadly sample the descriptor space while also reflecting practical considerations of fabrication accessibility and turnaround. Experimental outcomes for Al, Ta, TiW3, Pd, Pt, and Au are indicated by green or red circles outlining the data points, where green and red denote whether niobium oxide formation was observed to be suppressed or not. These data confirmed the broad trends predicted by the descriptor space, with successful barrier metals appearing toward the upper left and unsuccessful metals toward the lower right. We then used this data to train the first iteration of our logistic regression model as described above in Logistic regression, initiating the theory-experiment loop depicted in Figure 1. The predicted probabilities are encoded in the figure as the color of the inner solid circles, ranging from green ($P=0$) to red ($P=1$).
Metals were then prioritized for further testing based on their predicted probability of success. After each new set of experiments, the results were fed back into the model, which was retrained to improve its predictive accuracy. This created a closed theory-experiment loop that progressed over several iterations. The materials search space encompassed 27 candidate metals relevant to niobium-based junctions, selected to span transition-metal rows, d-band fillings, and selected s- and p-block native-oxide chemistries. Aluminum oxide provides the canonical barrier reference, while Be-, Zn-, and Bi-based oxides broadened the oxidation search space. By the most recent iteration, shown in Figure 4b, the model had successfully predicted the behavior of five new potential barrier materials: Zr, Hf, Ti, Mo, and Ir, with one unsuccessful case: W. The final model shows strong consistency with experiment and captures smooth, interpretable probability gradients across the descriptor space.
@@
To test whether the per oxygen oxide formation energy is sufficient to explain our logistic regression predictions, we next project a line of constant oxide formation energy (determined by logistical regression holding the slope of the line fixed) back onto the original descriptor space, as seen in Figure 4b. The resulting line separates the data just as well as the decision boundary learned by our original logistic regression model, indicating that per-oxygen oxide formation energy alone is sufficient to reproduce the model’s classification performance.
Figure 6: Per-oxygen oxide formation energy of various experimentally tested oxides, colored according to their success (green) or failure (red) to prevent oxidation of the niobium layer.
@@
Figure 7: Oxide formation energy versus lattice mismatch percentage against niobium: Metals have been colored according to their logistic predictions, along with outer circles indicating experimental results denoted by their color.
@@
Conclusion. This work establishes a closed-loop framework integrating defect-level energetics, logistic-regression modeling, and targeted experiments to guide functional materials design. Applied to superconducting qubit interfaces, it identifies and experimentally validates promising oxygen-barrier materials, including Zr, Ta, and Hf-based capping layers, while highlighting additional candidates such as Sc. Lattice matching further emerges as an important design criterion that refines materials selection, reducing the relative appeal of Hf while elevating Sc despite its cost. More broadly, the approach is applicable wherever physically meaningful, computable descriptors can be defined, enabling accelerated discovery of next-generation materials systems, particularly in experimentally data-sparse regimes limited by the pace of materials synthesis and characterization.
Acknowledgments. This work was supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0706. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force. This work made use of the Cornell Center for Materials Research shared instrumentation facility for the XPS analysis system. C. Méndez was supported by the US National Science Foundation under award PHY1549132, the Center for Bright Beams.
Data availability. The computational data underlying this study — including first-principles inputs and outputs, processed defect energetics, logistic-regression code, and links to the software used to generate the data — are publicly available on GitHub [50]. Experimental XPS measurements used to classify diffusion-barrier performance are available on Zenodo [51].